# Weak-Lensing Shear-Selected Galaxy Clusters from the Hyper Suprime-Cam Subaru Strategic Program: I. Cluster Catalog, Selection Function and Mass–Observable Relation

### v1

@@ Abstract
We present the first step towards deriving cosmological constraints through the abundances of galaxy clusters selected in a $510\,\mathrm{deg}^{2}$ weak-lensing aperture mass map, constructed with the Year-Three shear catalog from the Hyper Suprime-Cam Subaru Strategic Program. We adopt a conservative source galaxy selection to construct a sample of $129$ weak-lensing peaks with a signal-to-noise ratio above $4.7$. We use semi-analytical injection simulations to derive the selection function and the mass–observable relation of our sample. These results take into account complicated uncertainties associated with weak-lensing measurements, such as the non-uniform survey depth and the complex survey geometry, projection effects from uncorrelated large-scale structures, and the intrinsic alignment of source galaxies. We also propose a novel modeling framework to make parts of the mass–observable relation insensitive to assumed cosmological parameters. Such a framework not only offers a great computational advantage to cosmological studies, but can also benefit future astrophysical studies using shear-selected clusters. Our results are an important step towards utilizing these cluster samples that are constructed nearly independent of any baryonic assumptions in upcoming deep-and-wide lensing surveys from the Vera Rubin Observatory, Euclid, and the Nancy Grace Roman Space Telescope.
@@ 1 Introduction
In particular, observations utilizing the effect of weak gravitational lensing (hereinafter weak lensing; see Mandelbaum, 2018, for a comprehensive review) offer some of the strongest constraints on $S_{8}$ from the late universe. These observations measure small but coherent shape distortions, often referred to as shears, of distant galaxies (source) by intervening foreground structures (lens), giving us a direct probe of matter density fluctuations. The first cosmological application of weak lensing comes from the detection of cosmic shear variance (Bacon et al., 2000; Kaiser et al., 2000; Van Waerbeke et al., 2000; Wittman et al., 2000), and subsequent cosmological constraints from cosmic shear power spectra and correlation functions have shown the power of weak lensing with just the Gaussian summary statistics (Kilbinger, 2015). Meanwhile, it is of increasing interest to extract additional cosmological information from higher-order statistics in weak-lensing observations. Examples of these summary statistics include number counts of peaks (Dietrich & Hartlap, 2010; Hamana et al., 2015; Liu et al., 2015b, a; Kacprzak et al., 2016; Shan et al., 2018; Martinet et al., 2018; Harnois-Déraps et al., 2021; Zürcher et al., 2022; Liu et al., 2023; Marques et al., 2023) or higher moments from weak-lensing mass maps (Van Waerbeke et al., 2013; Petri et al., 2015; Peel et al., 2018; Chang et al., 2018; Gatti et al., 2020, 2022; Anbajagane et al., 2023), three-point correlation functions or bispectra (Takada & Jain, 2003, 2004; Dodelson & Zhang, 2005; Vafaei et al., 2010; Bergé et al., 2010; Semboloni et al., 2011; Fu et al., 2014), Minkowski functionals (Kratochvil et al., 2012; Petri et al., 2013; Vicinanza et al., 2019; Parroni et al., 2020), density split statistics (Gruen et al., 2018; Friedrich et al., 2018), and direct field-level inference (Jeffrey et al., 2020; Fluri et al., 2022). Obtaining accurate and competitive cosmological constraints from these higher-order statistics will serve as an important consistency test among weak-lensing probes and help break parameter degeneracies.
@@
In a series of two papers (this work and Chiu et al., 2024), we focus on constraining cosmology with high signal-to-noise ratio peaks detected on the weak-lensing aperture mass maps (Schneider, 1996; Oguri et al., 2021). These peaks are associated with massive dark matter halos where clusters of galaxies reside. The number counts of galaxy clusters, modeled by the halo mass function, are sensitive to both the geometry and the structure formation history in our universe (e.g., Allen et al., 2011). Tight cosmological constraints have been obtained with cluster samples selected with X-ray signals (Vikhlinin et al., 2009; Mantz et al., 2010, 2014; Chiu et al., 2023; Ghirardini et al., 2024) from the thermal bremsstrahlung emission, at millimeter wavelengths (Planck Collaboration et al., 2016; de Haan et al., 2016; Bocquet et al., 2019) by the thermal Sunyaev–Zel’dovich effect (Sunyaev & Zeldovich, 1972), and in the optical (Rozo et al., 2010; Abbott et al., 2020; Costanzi et al., 2021; Lesci et al., 2022a, b; Sunayama et al., 2023) via overdensities of galaxies. However, all of these samples rely on baryonic tracers of clusters and require complex astrophysical modeling to account for the selection effect and the mass–observable relation. Incorrect or insufficient modeling of these baryonic effects could lead to biases in the cosmological constraints (Salvati et al., 2020; Grandis et al., 2021). Cluster samples constructed from weak-lensing maps (Wittman et al., 2001; Schirmer et al., 2007; Miyazaki et al., 2007; Shan et al., 2012; Miyazaki et al., 2015, 2018; Oguri et al., 2021), on the other hand, allow us to determine the selection effect based purely on the theory of gravity, offering a direct link between the weak-lensing observable and the underlying halo mass. Shear-selected clusters, therefore, provide a powerful cosmology probe that complements both cosmic shear power spectrum and traditional cluster cosmology.
@@
Despite the direct link between the weak-lensing observable and the underlying projected density field, it is still essential to construct a comprehensive modeling framework that takes into account all physical and systematic effects associated with weak-lensing observations. In the past, weak-lensing peak counts were often modeled analytically by using halo models on a Gaussian random field (Fan et al., 2010; Shan et al., 2018) or semi-analytically by injecting synthetic halo profiles into N-body simulations (Lin & Kilbinger, 2015). Important systematic effects such as dilution by cluster member galaxies (Medezinski et al., 2018; Hamana et al., 2020; Oguri et al., 2021), intrinsic alignment of source galaxies (Kacprzak et al., 2016; Harnois-Déraps et al., 2021; Zhang et al., 2022), and other baryon physics (Osato et al., 2015; Weiss et al., 2019; Coulton et al., 2020; Osato et al., 2021; Lee et al., 2023) are hard to model and can only be accounted for using nuisance parameters with strong priors. On the other hand, ray-tracing simulations have also been carried out to study weak-lensing peak counts (Dietrich & Hartlap, 2010; Liu et al., 2015a; Kacprzak et al., 2016). However, these simulations are generally expensive to compute and can only be carried out on a limited grid of cosmological parameters.
@@ 2.1 Weak-lensing shape catalog
The shape catalog from the HSC-SSP S19A internal data release (Li et al., 2022, hereinafter the HSC-Y3 data) contains roughly 36 million galaxies across six fields (XMM, GAMA09H, WIDE12H, GAMA15H, VVDS, HECTOMAP) [^1] The complete HSC survey consists of only three distinct patches. However, as of year three, the survey footprint is still fragmented. with an averaged $19.9~{}\mathrm{arcmin}^{-2}$ effective galaxy number density. The galaxy shapes are measured with the re-Gaussianization method (Hirata & Seljak, 2003) and the shear estimation bias is calibrated using image simulations (Mandelbaum et al., 2018a). A series of null tests performed in Li et al. (2022) has demonstrated that various residual systematics are well under controlled, making this a catalog suitable for accurate weak-lensing measurements.
@@
Here, the probability density function $P(z)$ for each source galaxy is obtained with an empirical fitting method from the Direct Empirical Photometric code (DEmP; Hsieh & Yee, 2014; Tanaka et al., 2018; Nishizawa et al., 2020). With this conservative cut, the total number of source galaxies reduces to roughly 16.9 million galaxies with an averaged $9.4\,\mathrm{arcmin}^{-2}$ effective galaxy number density. The combined probability density function of source galaxies’ redshift before and after this selection is shown as the dashed gray and solid green curve in the left panel of Fig. 1.
@@
Figure 1: Left: Combined probability density function of source galaxies’ photometric redshift in the HSC-Y3 shape catalog before (dashed grey) and after (solid green) the selection criterion given in eq. (1). The redshift distribution of weak-lensing peaks with signal-to-noise ratio larger than 4.7 is shown in the blue histogram. The redshift of the weak-lensing peaks is obtained from cross-matching with optical cluster catalogs as described in Sec. 2.3. Right: The aperture mass filters for convergence (red) and tangential shear (blue) adopted in this work.
@@
Figure 2: Mass maps derived from the HSC-Y3 data with the truncated isothermal filter given in Fig. 1. Weak-lensing peaks with a signal-to-noise ratio larger than 4.7 are circled in green. These peaks are matched to optical clusters in the following order: CAMIRA cluster with a high $P(N_{\mathrm{mem}},z|\nu)$ (purple square); low-redshift clusters found in the SDSS (brown hexagon); any other optical clusters with $N_{\mathrm{mem}}\geq 20$ (blue diamond).
@@ 2.2 Construction of Mass Maps
In this work, we adopt a truncated isothermal profile (Schneider, 1996) {widetext}
$U(\theta)=\begin{cases}1&\left(\theta\leq x_{1}\theta_{R}\right)\\ \displaystyle\frac{1}{1-c}\left(\frac{x_{1}\theta_{R}}{\sqrt{\left(\theta-x_{1}\theta_{R}\right)^{2}+\left(x_{1}\theta_{R}\right)^{2}}}-c\right)&\left(x_{1}\theta_{R}\leq\theta\leq x_{2}\theta_{R}\right)\\ \displaystyle\frac{b}{\theta_{R}^{3}}\left(\theta_{R}-\theta\right)^{2}\left(\theta-\alpha\theta_{R}\right)&\left(x_{2}\theta_{R}\leq\theta\leq\theta_{R}\right)\\ 0&\left(\theta_{R}\leq\theta\right)\end{cases}$ (6)
where the constants $\alpha,b,c$ are chosen so that the filter $U$ and its first derivative are both continuous at $\theta=x_{2}\theta_{R}$ and the normalization condition in Eq. (3) is satisfied. The overall shape of such a filter is designed to roughly match the surface density distribution of a galaxy cluster. Meanwhile, a constant filter for the convergence at $\theta\leq x_{1}\theta_{R}$ makes us insensitive to the potential non-linear and baryonic systematic uncertainties in the innermost region of a cluster, and a smoothly truncated filter reduces possible discontinuities in our aperture mass maps from the complex survey geometry. Following Oguri et al. (2021), we adopt $x_{1}=0.121$, $x_{2}=0.36$, and $\theta_{R}=16.6\,\mathrm{arcmin}$, which has been shown to be the most effective in mitigating the dilution effect from member galaxies and in suppressing the noise from uncorrelated large-scale structures. The resulting filter $U(\theta)$ and $Q(\theta)$ is shown in the right panel of Fig. 1.
@@
Our weak-lensing maps are sampled on a two-dimensional rectangular grid in the $\boldsymbol{\mathbf{\theta}}$-coordinate with a pixel size of $0.25$ arcmin. Fig. 2 shows six signal-to-noise maps derived from the truncated isothermal filter. For detailed procedure of calculating these weak-lensing maps, we refer the reader to Sec. 3.4 of Oguri et al. (2021).
@@ 2.3 Cluster catalogs
For each weak-lensing peaks, we search for optical clusters within a $10$ arcmin radius and keep clusters that are within $1.5\,h^{-1}\mathrm{Mpc}$ using the redshift of the matched optical clusters. As each weak-lensing peak could be matched to multiple optical clusters, we develop the following framework to determine the best-matched optical counterpart. First, we search for potential counterparts within the CAMIRA catalog using the above-mentioned criteria based on both angular and physical distances. For each of the cross-matched CAMIRA candidate of the observed weak-lensing peak with a signal-to-noise ratio $\nu$, we adopt the richness-to-mass relation from Murata et al. (2019) to estimate the probability of observing the counterpart with a richness of $N_{\mathrm{mem}}$ at redshift $z_{\mathrm{cl}}$, denoting as $P(N_{\mathrm{mem}},z_{\mathrm{cl}}|\nu)$ (See Appendix A for more details). We define a cross-matched CAMIRA candidate as a high-confidence counterpart if the richness and redshift of that cluster fall within the $3\sigma$ confidence interval around the peak of the distribution $P(N_{\mathrm{mem}},z_{\mathrm{cl}}|\nu)$. The best-matched optical counterpart is chosen as the most likely among all the high-confidence counterparts. If a peak does not admit a high-confidence counterpart, we then search for the nearest low-$z$ counterpart in the redMaPPer and WHL catalogs that has $z\leq 0.2$ and $N_{\mathrm{mem}}\geq 20$. Lastly, a peak is said to have any optical counterpart if it is matched to any optical clusters with $N_{\mathrm{mem}}\geq 15$ in either CAMIRA, redMaPPer or WHL cluster catalogs.
@@
Fig. 2 shows an example of a weak-lensing peak sample constructed with $\nu_{\mathrm{min}}=4.7$. Under this signal-to-noise ratio threshold, there are in total $129$ peaks identified in the entire HSC-Y3 footprint. Among which, $112$ are found to have high-confidence optical counterparts within the CAMIRA catalog; $2$ peaks do not have high-confidence CAMIRA matches but admit low-$z$ counterparts in SDSS, and $12$ are matched to optical counterparts with $N_{\mathrm{mem}}\geq 15$ and are within $1.5~{}h^{-1}\mathrm{Mpc}$ to the peak. While there are $3$ peaks that do not have any optical counterpart under the distance and richness criteria given above, two of them (located at $12\mathrm{h}39\mathrm{m}47.0645\mathrm{s},-00\mathrm{d}33\mathrm{m}52.8076\mathrm{s}$ and $22\mathrm{h}24\mathrm{m}51.0696\mathrm{s}+03\mathrm{d}52\mathrm{m}47.9063\mathrm{s}$) are surrounded by multiple lower-richness CAMIRA clusters and the remaining one ($02\mathrm{h}33\mathrm{m}40.6278\mathrm{s}-03\mathrm{d}56\mathrm{m}00.1856\mathrm{s}$) is near optical clusters identified by the Canada-France-Hawaii Telescope at $z\sim 0.5$ (Durret et al., 2011; Sarron et al., 2018). Therefore, these peaks are likely to be the results of chance projection of multiple halos along the line-of-sight. Such a scenario is taken into account in our modeling framework given in Sec. 3. A list of weak-lensing peaks in the HSC-Y3 footprint with $\nu\geq 4.7$ and their best-mateched optical counterpart is given in Table 3.
The redshift distribution of the best-matched optical counterpart for weak-lensing peaks with $\nu\geq 4.7$ is given as the blue histogram in the left panel of Fig. 1. We see that all but one weak-lensing peaks are located at $z\leq 0.7$. As we only utilize source galaxies with significant probability of being above redshift $0.7$, combining with the fact that we exclude signal from the innermost region of the halo, the weak-lensing observable we obtained is safe from the dilution of cluster member galaxies.
Fig. 3 shows the number of peaks as a function of the selection threshold $\nu_{\mathrm{min}}$ and the percentage of peaks with (high confidence) optical counterparts. We observe that $96\%\,(100\%)$ of peaks with $\nu\geq 4.0\,(5.0)$ admit optical counterpart, indicating these weak-lensing peaks to be an effective way to select galaxy clusters without any baryonic assumptions. In Part II of our analysis (Chiu et al., 2024), we will utilize these peak samples constructed with different selection threshold $\nu_{\mathrm{min}}$ to test the robustness of our cosmology constraints.
@@
Table 3 is published in its entirety in the machine-readable format [^2] https://github.com/inonchiu/hsc_shear_selected_clusters. A portion is shown here for guidance regarding its form and content.

### v2

@@ Abstract
We present the first step toward deriving cosmological constraints through the abundances of galaxy clusters selected in a $510\,\mathrm{deg}^{2}$ weak-lensing aperture mass map, constructed with the Year-Three shear catalog from the Hyper Suprime-Cam Subaru Strategic Program. We adopt a conservative source galaxy selection to construct a sample of $129$ weak-lensing peaks with a signal-to-noise ratio above $4.7$. We use semi-analytical injection simulations to derive the selection function and the mass–observable relation of our sample. These results take into account complicated uncertainties associated with weak-lensing measurements, such as the non-uniform survey depth and the complex survey geometry, projection effects from uncorrelated large-scale structures, and the intrinsic alignment of source galaxies. We also propose a novel modeling framework to make parts of the mass–observable relation insensitive to assumed cosmological parameters. Such a framework not only offers a great computational advantage to cosmological studies, but can also benefit future astrophysical studies using shear-selected clusters. Our results are an important step toward utilizing these cluster samples that are constructed nearly independent of any baryonic assumptions in upcoming deep-and-wide lensing surveys from the Vera Rubin Observatory, Euclid, and the Nancy Grace Roman Space Telescope.
@@ 1 Introduction
In particular, observations utilizing the effect of weak gravitational lensing (hereinafter weak lensing; see Mandelbaum, 2018, for a comprehensive review) offer some of the strongest constraints on $S_{8}$ from the late universe. These observations measure small but coherent shape distortions, often referred to as shears, of distant galaxies (source) by intervening foreground structures (lens), giving us a direct probe of matter density fluctuations. The first cosmological application of weak lensing comes from the detection of cosmic shear variance (Bacon et al., 2000; Kaiser et al., 2000; Van Waerbeke et al., 2000; Wittman et al., 2000), and subsequent cosmological constraints from cosmic shear power spectra and correlation functions have shown the power of weak lensing with just the Gaussian summary statistics (Kilbinger, 2015). Meanwhile, it is of increasing interest to extract additional cosmological information from higher-order statistics in weak-lensing observations. Examples of these summary statistics include number counts of peaks (Dietrich & Hartlap, 2010; Hamana et al., 2015; Liu et al., 2015b, a; Kacprzak et al., 2016; Shan et al., 2018; Martinet et al., 2018; Harnois-Déraps et al., 2021; Zürcher et al., 2022; Liu et al., 2023; Marques et al., 2024) or higher moments from weak-lensing mass maps (Van Waerbeke et al., 2013; Petri et al., 2015; Peel et al., 2018; Chang et al., 2018; Gatti et al., 2020, 2022; Anbajagane et al., 2023), three-point correlation functions or bispectra (Takada & Jain, 2003, 2004; Dodelson & Zhang, 2005; Vafaei et al., 2010; Bergé et al., 2010; Semboloni et al., 2011; Fu et al., 2014), Minkowski functionals (Kratochvil et al., 2012; Petri et al., 2013; Vicinanza et al., 2019; Parroni et al., 2020), density split statistics (Gruen et al., 2018; Friedrich et al., 2018), and direct field-level inference (Jeffrey et al., 2020; Fluri et al., 2022). Obtaining accurate and competitive cosmological constraints from these higher-order statistics will serve as an important consistency test among weak-lensing probes and help break parameter degeneracies.
@@
In a series of two papers (this work and Chiu et al., 2024), we focus on constraining cosmology with high signal-to-noise ratio peaks detected on weak-lensing aperture mass maps (Schneider, 1996; Oguri et al., 2021). These peaks are associated with massive dark matter halos where clusters of galaxies reside. The number counts of galaxy clusters, modeled by the halo mass function, are sensitive to both the geometry and the structure formation history in our universe (e.g., Allen et al., 2011). Tight cosmological constraints have been obtained with cluster samples selected with X-ray signals (Vikhlinin et al., 2009; Mantz et al., 2010, 2014; Chiu et al., 2023; Ghirardini et al., 2024) from the thermal bremsstrahlung emission, at millimeter wavelengths (Planck Collaboration et al., 2016; de Haan et al., 2016; Bocquet et al., 2019) by the thermal Sunyaev–Zel’dovich effect (Sunyaev & Zeldovich, 1972), and in the optical (Rozo et al., 2010; Abbott et al., 2020; Costanzi et al., 2021; Lesci et al., 2022a, b; Sunayama et al., 2024) via overdensities of galaxies. However, all of these samples rely on baryonic tracers of clusters and require complex astrophysical modeling to account for the selection effect and the mass–observable relation. Incorrect or insufficient modeling of these baryonic effects could lead to biases in the cosmological constraints (Salvati et al., 2020; Grandis et al., 2021). Cluster samples constructed from weak-lensing maps (Wittman et al., 2001; Schirmer et al., 2007; Miyazaki et al., 2007; Shan et al., 2012; Miyazaki et al., 2015, 2018; Oguri et al., 2021), on the other hand, allow us to determine the selection effect based purely on the theory of gravity, offering a direct link between the weak-lensing observable and the underlying halo mass. Shear-selected clusters, therefore, provide a powerful cosmology probe that complements both the cosmic shear power spectrum and traditional cluster cosmology.
@@
Despite the direct link between the weak-lensing observable and the underlying projected density field, it is still essential to construct a comprehensive modeling framework that takes into account all physical and systematic effects associated with weak-lensing observations. In the past, weak-lensing peak counts were often modeled analytically by using halo models on a Gaussian random field (Fan et al., 2010; Shan et al., 2018) or semi-analytically by injecting synthetic halo profiles into N-body simulations (Lin & Kilbinger, 2015). Important systematic effects such as dilution by cluster member galaxies (Medezinski et al., 2018; Hamana et al., 2020; Oguri et al., 2021), intrinsic alignment of source galaxies (Kacprzak et al., 2016; Harnois-Déraps et al., 2021; Zhang et al., 2022), and other baryon physics (Osato et al., 2015; Weiss et al., 2019; Coulton et al., 2020; Osato et al., 2021; Lee et al., 2023; Broxterman et al., 2024) are hard to model and can only be accounted for using nuisance parameters with strong priors. On the other hand, ray-tracing simulations have also been carried out to study weak-lensing peak counts (Dietrich & Hartlap, 2010; Liu et al., 2015a; Kacprzak et al., 2016). However, these simulations are generally expensive to compute and can only be carried out on a limited grid of cosmological parameters.
@@ 2.1 Weak-lensing shape catalog
The shape catalog from the HSC-SSP S19A internal data release (Li et al., 2022, hereinafter the HSC-Y3 data) contains roughly 36 million galaxies across six fields (XMM, GAMA09H, WIDE12H, GAMA15H, VVDS, HECTOMAP) [^1] The complete HSC survey consists of only three distinct patches. However, as of year three, the survey footprint is still fragmented. with an averaged $19.9~{}\mathrm{arcmin}^{-2}$ effective galaxy number density. The galaxy shapes are measured with the re-Gaussianization method (Hirata & Seljak, 2003) and the shear estimation bias is calibrated using image simulations (Mandelbaum et al., 2018a). The mean shear values, the stacked cross shear signals around galaxies, and the star-galaxy cross-correlation functions of the Y3 catalog have been demonstrated to be mostly consistent with zero (Li et al., 2022), enabling accurate weak-lensing measurements.
@@
Here, the probability density function $P(z)$ for each source galaxy is obtained with an empirical fitting method from the Direct Empirical Photometric code (DEmP; Hsieh & Yee, 2014; Tanaka et al., 2018; Nishizawa et al., 2020). With this conservative cut, the total number of source galaxies reduces to roughly 16.9 million galaxies with an averaged $9.4\,\mathrm{arcmin}^{-2}$ effective galaxy number density. The combined probability density function of source galaxies’ redshift before and after this selection is shown as the dashed gray and solid green curve in the left panel of Figure 1.
@@
Figure 1: Left: probability density function of source galaxies’ photometric redshift in the HSC-Y3 shape catalog before (dashed grey) and after (solid green) the selection criterion given in Equation (1). The redshift distribution of weak-lensing peaks with signal-to-noise ratio larger than 4.7 is shown in the blue histogram. The redshift of the weak-lensing peaks is obtained from cross-matching with optical cluster catalogs as described in Sec. 2.3. Right: aperture mass filters for convergence (red) and tangential shear (blue) adopted in this work.
@@
Figure 2: Mass maps derived from the HSC-Y3 data with the truncated isothermal filter given in Figure 1. Weak-lensing peaks with a signal-to-noise ratio larger than 4.7 are circled in green. These peaks are matched to optical clusters in the following order: CAMIRA cluster with a high $P(N_{\mathrm{mem}},z|\nu)$ (purple square); low-redshift clusters found in the SDSS (brown hexagon); any other optical clusters with richness $N_{\mathrm{mem}}\geq 20$ (blue diamond).
@@ 2.2 Construction of Mass Maps
In this work, we adopt a truncated isothermal profile (Schneider, 1996)
$U(\theta)=$ (6)
$\begin{cases}1&\left(\theta\leq x_{1}\theta_{R}\right)\\ \displaystyle\frac{1}{1-c}\left(\frac{x_{1}\theta_{R}}{\sqrt{\left(\theta-x_{1}\theta_{R}\right)^{2}+\left(x_{1}\theta_{R}\right)^{2}}}-c\right)&\left(x_{1}\theta_{R}\leq\theta\leq x_{2}\theta_{R}\right)\\ \displaystyle\frac{b}{\theta_{R}^{3}}\left(\theta_{R}-\theta\right)^{2}\left(\theta-\alpha\theta_{R}\right)&\left(x_{2}\theta_{R}\leq\theta\leq\theta_{R}\right)\\ 0&\left(\theta_{R}\leq\theta\right)\end{cases}$
where the constants $\alpha,b,c$ are chosen so that the filter $U$ and its first derivative are both continuous at $\theta=x_{2}\theta_{R}$ and the normalization condition in Equation (3) is satisfied. The overall shape of such a filter is designed to roughly match the surface density distribution of a galaxy cluster. Meanwhile, a constant filter for the convergence at $\theta\leq x_{1}\theta_{R}$ makes us insensitive to the potential non-linear and baryonic systematic uncertainties in the innermost region of a cluster, and a smoothly truncated filter reduces possible discontinuities in our aperture mass maps from the complex survey geometry. Following Oguri et al. (2021), we adopt $x_{1}=0.121$, $x_{2}=0.36$, and $\theta_{R}=16.6\,\mathrm{arcmin}$, which has been shown to be the most effective in mitigating the dilution effect from member galaxies and in suppressing the noise from uncorrelated large-scale structures. The resulting filter $U(\theta)$ and $Q(\theta)$ is shown in the right panel of Figure 1.
@@
Our weak-lensing maps are sampled on a two-dimensional rectangular grid in the $\boldsymbol{\mathbf{\theta}}$-coordinate with a pixel size of $0.25$ arcmin. Figure 2 shows six signal-to-noise maps derived from the truncated isothermal filter. For detailed procedure of calculating these weak-lensing maps, we refer the reader to Sec. 3.4 of Oguri et al. (2021).
@@ 2.3 Cluster catalogs
For each weak-lensing peaks, we search for optical clusters within a $10$ arcmin radius and keep clusters that are within $1.5\,h^{-1}\mathrm{Mpc}$ using the redshift of the matched optical clusters. As each weak-lensing peak could be matched to multiple optical clusters, we develop the following framework to determine the best-matched optical counterpart. First, we search for potential counterparts within the CAMIRA catalog using the above-mentioned criteria based on both angular and physical distances. For each of the cross-matched CAMIRA candidate of the observed weak-lensing peak with a signal-to-noise ratio $\nu$, we adopt the richness-to-mass relation from Murata et al. (2019) to estimate the probability of observing the counterpart with a richness of $N_{\mathrm{mem}}$ at redshift $z_{\mathrm{cl}}$, denoting as $P(N_{\mathrm{mem}},z_{\mathrm{cl}}|\nu)$ (See Appendix A for more details). Here, the richness parameter $N_{\mathrm{mem}}$ is an estimate of the number of cluster member galaxies and is derived from a weighted sum of background-subtracted red galaxy number density around the brightest cluster galaxy (Oguri, 2014; Oguri et al., 2018). We define a cross-matched CAMIRA candidate as a high-confidence counterpart if the richness and redshift of that cluster fall within the $3\sigma$ confidence interval around the peak of the distribution $P(N_{\mathrm{mem}},z_{\mathrm{cl}}|\nu)$. The best-matched optical counterpart is chosen as the most likely among all the high-confidence counterparts. If a peak does not admit a high-confidence counterpart, we then search for the nearest low-$z$ counterpart in the redMaPPer and WHL catalogs that has $z\leq 0.2$ and $N_{\mathrm{mem}}\geq 20$. Lastly, a peak is said to have any optical counterpart if it is matched to any optical clusters with $N_{\mathrm{mem}}\geq 15$ in either CAMIRA, redMaPPer or WHL cluster catalogs  [^2] We note that while the definition of richness is different between the three optical catalogs, here we adopt a common selection threshold for simplicity. For the CAMIRA catalog, $N_{\mathrm{mem}}\geq 15$ roughly corresponds to clusters with $M_{\mathrm{200m}}\geq 10^{14}~{}h^{-1}M_{\odot}$ (Oguri et al., 2018). .
@@
Figure 2 shows an example of a weak-lensing peak sample constructed with $\nu_{\mathrm{min}}=4.7$. Under this signal-to-noise ratio threshold, there are in total $129$ peaks identified in the entire HSC-Y3 footprint. Among which, $112$ are found to have high-confidence optical counterparts within the CAMIRA catalog; $2$ peaks do not have high-confidence CAMIRA matches but admit low-$z$ counterparts in SDSS, and $12$ are matched to optical counterparts with $N_{\mathrm{mem}}\geq 15$ and are within $1.5~{}h^{-1}\mathrm{Mpc}$ to the peak. While there are $3$ peaks that do not have any optical counterpart under the distance and richness criteria given above, two of them (located at $12\mathrm{h}39\mathrm{m}47.0645\mathrm{s},-00\mathrm{d}33\mathrm{m}52.8076\mathrm{s}$ and $22\mathrm{h}24\mathrm{m}51.0696\mathrm{s}+03\mathrm{d}52\mathrm{m}47.9063\mathrm{s}$) are surrounded by multiple lower-richness CAMIRA clusters and the remaining one ($02\mathrm{h}33\mathrm{m}40.6278\mathrm{s}-03\mathrm{d}56\mathrm{m}00.1856\mathrm{s}$) is near optical clusters identified by the Canada-France-Hawaii Telescope at $z\sim 0.5$ (Durret et al., 2011; Sarron et al., 2018). Therefore, these peaks are likely to be the results of chance projection of multiple halos along the line-of-sight. Such a scenario is taken into account in our modeling framework given in Sec. 3. A list of weak-lensing peaks in the HSC-Y3 footprint with $\nu\geq 4.7$ and their best-mateched optical counterpart is given in Figure 3.
The redshift distribution of the best-matched optical counterpart for weak-lensing peaks with $\nu\geq 4.7$ is given as the blue histogram in the left panel of Figure 1. We see that all but one weak-lensing peaks are located at $z\leq 0.7$. As we only utilize source galaxies with significant probability of being above redshift $0.7$, combining with the fact that we exclude signal from the innermost region of the halo, the weak-lensing observable we obtained is safe from the dilution of cluster member galaxies.
Figure 3 shows the number of peaks as a function of the selection threshold $\nu_{\mathrm{min}}$ and the percentage of peaks with (high confidence) optical counterparts. We observe that $96\%\,(100\%)$ of peaks with $\nu\geq 4.0\,(5.0)$ admit optical counterpart, indicating these weak-lensing peaks to be an effective way to select galaxy clusters without any baryonic assumptions. Following Miyazaki et al. (2018) and Oguri et al. (2021), in this work, we focus on peaks with $\nu\geq 4.7$. In Part II of our analysis (Chiu et al., 2024), we will utilize peak samples constructed with different selection threshold $\nu_{\mathrm{min}}$ to test the robustness of our cosmological constraints.
@@
Figure 3 is published in its entirety in the machine-readable format [^3] https://github.com/inonchiu/hsc_shear_selected_clusters. A portion is shown here for guidance regarding its form and content.